OpenStudy (anonymous):

Perform the modular arithmetic operation. [(5 - 9) ∙ 3] (mod 5) A. 1 B. 2 C. 0 D. 3

OpenStudy (anonymous):

a=qn + r -12=-2(5)+r -12+10=-2 r=-2 Not sure if this is correct...it's not one of your choices either.

OpenStudy (anonymous):

yeah I'm probably incorrect here. I'm trying to learn "modular arithmetic" on the fly here (never heard of it before)...